OpenStudy (amillen16):

Lie is solving the quadratic equation by completing the square. 3x2 + 18x – 8 = 0 3x2 + 18x = 8 what should he do next?

OpenStudy (zzr0ck3r):

\(ax^2+bx+c=0\iff ax^2+bx=-c\iff a(x^2+\frac{b}{a}x)=-c\)

OpenStudy (amillen16):

thanks